OpenStudy (calculusxy):

Suzanne bought a sweater at the sale price of $25. The original cost of the sweater was $40. What percent represents the discount that Suzanne received when buying the sweater?

OpenStudy (calculusxy):

@hartnn @TuringTest

OpenStudy (calculusxy):

15/4000?

OpenStudy (perl):

$$ \frac{|new price - old price | }{old price} \cdot 100 $$

OpenStudy (calculusxy):

What would I do if i have the equation: Percent x whole = part?

OpenStudy (perl):

should be 15/40 * 100 = 1500 / 40